OpenStudy (anonymous):

Find the product of each complex number and its conjugate. 20. 5 23. 9i

OpenStudy (anonymous):

\[9i\] as in imaginary number

zepdrix (zepdrix):

\[\Large\rm 5=5+0i\] \[\Large\rm (5+0i)(5-0i)=?\]

OpenStudy (anonymous):

So if a term is missing, either real number or imaginary, it is an unwritten zero?

zepdrix (zepdrix):

yes! :)

zepdrix (zepdrix):

Certainly makes your multiplication easier huh? ^^

zepdrix (zepdrix):

Careful with 23 though,\[\Large\rm (0+9i)(0-9i)\neq 9i\cdot 9i\]

OpenStudy (anonymous):

I've got it from here! 20. 5 = 25 23. 9i = 9

zepdrix (zepdrix):

9^2 for 23, ya? :)

OpenStudy (anonymous):

\[-9i^2 = -9 * -1 = 9 \] right?

zepdrix (zepdrix):

\[\Large\rm (0+9i)(0-9i)= 9i\cdot(-9i)=-(9i)^2=-9^2i^2\]

OpenStudy (anonymous):

Oh woops, just a careless mistake I made. Thanks!

zepdrix (zepdrix):

np c: